### v1

# Continual Imitation Learning for Prosthetic Limbs

Abstract
Lower limb amputations and neuromuscular impairments severely restrict mobility, necessitating advancements beyond conventional prosthetics. Motorized bionic limbs offer promise, but their utility depends on mimicking the evolving synergy of human movement in various settings. In this context, we present a novel model for bionic prostheses’ application that leverages camera-based motion capture and wearable sensor data, to learn the synergistic coupling of the lower limbs during human locomotion, empowering it to infer the kinematic behavior of a missing lower limb across varied tasks, such as climbing inclines and stairs. We propose a model that can multitask, adapt continually, anticipate movements, and refine. The core of our method lies in an approach which we call “multitask prospective rehearsal” that anticipates and synthesizes future movements based on the previous prediction and employs a corrective mechanism for subsequent predictions. We design an evolving architecture that merges lightweight, task-specific modules on a shared backbone, ensuring both specificity and scalability. We empirically validate our model against various baselines using real-world human gait datasets, including experiments with transtibial amputees, which encompass a broad spectrum of locomotion tasks. The results show that our approach consistently outperforms baseline models, particularly under scenarios affected by distributional shifts, adversarial perturbations, and noise.
I Introduction
Lower limb amputation and neuromuscular disorders detrimentally affect natural locomotion, compromising individuals’ quality of life [1]. Those afflicted often depend on assistive technologies like prosthetics or orthoses to regain daily mobility [1]. However, conventional passive prosthetics often fall short in replicating natural gait during diverse activities, from basic transitions like standing from a seated position to more dynamic actions like running or navigating slopes [1]. The advent of powered prosthetics, equipped with integrated motors, promises a more naturalistic gait. Nonetheless, this advancement requires a model capable of effectively approximating the complexities of human gait synergy to accurately estimate the necessary motor commands. While finite-state machines are adequate for rudimentary scenarios [2, 3], their construction becomes infeasible for finer gait phase resolution or multiple locomotion tasks [4]. A promising alternative involves leveraging models trained on able-bodied human demonstrations to intuitively infer amputee limb motion. This methodology not only replicates gait’s inherent fluidity but also facilitates seamless transitions across varying gaits, eliminating the need for rigid rules or heuristics. The learned gait models could be personalized by training on a wide variety of human demonstrations, such that references with similar anthropomorphic features as that of the physically impaired subject are included.
Figure 1: Overview of the multitask prospective rehearsal-based training: In the pretraining phase, the multi gait predictor model, $f$, is trained to forecast the desired joint profile ${y}_{i}$ from the history of sensor states $\{\mathbf{X}_{i-T}...\mathbf{X}_{i}\}$, while the prospective model $g$ aims to predict the subsequent sensor state $\mathbf{X}_{i+1}$ from $\mathbf{X}_{i}$ and the target joint profiles ${y}_{i}$. For prospective rehearsal, the multi gait model predictions $\hat{{y}}_{k}$ are fed along with the current sensor state $\mathbf{X}_{k}$ to the prospective model to project the next sensor state $\hat{\mathbf{X}}_{k+1}$. Multi gait predictor, $f$ is then refined with this projected state, $\hat{\mathbf{X}}_{k+1},k=1...|\mathcal{X}_{val}|-1$ against the actual future output $y_{k+1}\in\mathcal{Y}_{val}$, to compensate the effect of its prediction errors.
Recent studies attest to the efficacy of learning-based models in gait and prosthetic behavior estimation [5, 6, 7]. However, most of these works have focused on either a single locomotion task or jointly training multiple modes in the context of multi-locomotion scenarios. Further, prior works have framed learning from human demonstration as a standard supervised learning task, which assumes independence across time. This oversight can cascade predictive errors, deviating significantly from the original training distribution (refer also to prior work on imitation learning, such as [8, 9]). Moreover, the multifaceted dynamics of human gait, which can shift based on varying terrains, speeds, fatigue levels, or specific locomotion tasks, underscore the necessity for a model that fluidly adapts to these evolving conditions.
To achieve this objective, we present a multitask, continually adaptive gait synergy approximation model tailored for various locomotion tasks, capable of adapting to evolving gait patterns and refining itself by integrating the impact of prediction errors. Central to our method is what we call the “multitask prospective rehearsal” (Figure 1) that prepares the model to anticipate and handle potential trajectories that may arise from its predictions, creating a seamless connection between continual adaptation and the integration of prediction errors for model refinement. Unlike conventional models that often rely on retrospective analysis of movement, our method is designed to prospectively imagine and synthesize potential future locomotion patterns. By synthesizing future states, the model is, in essence, creating its own new ’unseen’ data to practice on, which could lead to better generalization when encountering actual new data. The model refines its parameters incrementally as it updates itself with this enhanced data, negating the need for a simulator or new training data.
We conduct a wide range of experiments in continual and joint learning settings with different model architectures and backbones on three real-world gait datasets, including our own patient (transtibial amputee) dataset, which we release for future use. We establish that our model outperforms many baseline models with multiple benchmarks.
Our study is the first to approach multi-gait adaptation in bionic prostheses as an error-aware multitask continual adaptation problem. Additionally, it is among the few that merge multitask adaptation with task-incremental learning in time series regression. Our methodology aims to produce set-point signals for bionic prostheses, achieving human-like movement fidelity and offering individuals with physical impairments the possibility of regaining near-normal functionality.
II Related Work
Learning-based Gait Behavior Models. In the field of predictive gait analytics, a variety of learning-based models have emerged, utilizing data from motion capture systems and wearable sensors to estimate a range of variables for diverse objectives. Ardestani et al. [10] modeled the medial knee contact force prediction using a feedforward artificial neural network (FFAN) based on motion capture data, offering a computational alternative to traditional techniques. Zhang et al. [11] harnessed wearable sensors combined with support vector regression to extract essential gait parameters efficiently, minimizing the need for extensive subject training and effectively addressing inter-subject variability. Concurrently, Rai et al. [12] employed recurrent neural networks (RNNs), integrating data from wearable sensors, to yield predictions of ankle angles from various ambulatory activities. Wang et al. [13] showcased the precision of a Gated Recurrent Unit (GRU) model for gait prediction from optical motion capture data. Livolsi et al. [14] introduced discrete wavelet transform methods for continuous gait phase extraction from hip encoder data. Kim et al. [15] reported the use of Long Short-Term Memory (LSTM) models to predict gait phases, analyzing short-term historical data on vertical loads and limb kinematics. The k-nearest neighbors (kNN) algorithm has been calibrated to work with pressure sensor inputs, enhancing the precision of walking phase prediction for prosthetic control systems [16]. To create more adaptive and responsive systems, Dey et al. [17] incorporated estimators that adjust to new training data within decision learner ensembles, catering to individual gait patterns. Waugh et al. [18] proposed a canonical dynamical system (CDS) model leveraging the Fourier series representation of gait. Yet, these solutions have generally been either constrained to single tasks, did not incorporate the effect of prediction errors, or have not tackled the problem of forgetting previously learned tasks when new ones are introduced.
Multitask Adaptive Learning. A principal challenge in adaptive multitask learning is the phenomenon of catastrophic forgetting [19, 20], where newly acquired knowledge can cause the loss of previously learned information. To combat this, different continual learning strategies have been developed. Regularization-based methods like Elastic Weight Consolidation (EWC) [21] and Synaptic Intelligence (SI) [22] introduce regularization terms to guide the model towards a parameter space optimized for low error across both previous and new tasks. Evolving architectural methods such as Progressive Neural Networks (PNNs) [23] expand the model’s structure by introducing new, task-specific parameters to sustain performance across tasks. Rehearsal or replay methods, including Experience Replay (ER) and Gradient Episodic Memory (GEM) [24], maintain a sample of previous tasks’ data to be revisited during new task training, reinforcing the model’s exposure to the old data distribution. Rehearsal approaches have empirically proved to be the most effective ones among several approaches developed [25]. However, these techniques have a retrospective nature and are meant for treating data samples independently. In the context of bionic prosthetics, where predictive models interact with a dynamic environment, anticipating the impact of model predictions on subsequent data inputs is critical.
Model-based RL Model-based RL strategies generally focus on developing a representation of the environment by gathering data from interactions driven by a specific policy. This constructed model of the environment facilitates a level of planning, enabling the estimation of potential outcomes for future states [26, 27, 28, 29, 30]. Typically, these methods rely on a reward signal and online interactions with the environment. With the advent of offline reinforcement [31, 32] and model-based imitation learning [33, 34], access to the reward signals or online interactions can be bypassed [35]. As a result, it becomes feasible to simulate interactions and learn from these simulations, eliminating the necessity for real-time data acquisition. While model-based RL uses the world model for planning the agent’s actions based on its outcomes, we use it to generate rehearsal samples to fine-tune the prediction model to correct its own errors. Another major difference is that we use the world model in a multitask learning setting, whereas most of the model-based RL deals with single-task settings.
III Method
Let $\mathbf{X}_{t}=\mathbf{x}_{t,1},\ldots,\mathbf{x}_{t,N_{t}}\in\mathbb{R}^{d}$ be a time series of sensor states (3D body and joint angles, angular velocities, and linear accelerations) during a locomotion task $t$ (such as level ground walking, ascending a ramp, or climbing stairs) and let $\mathbf{y}_{t}=y_{t,1},\ldots,y_{t,N_{t}}\in\mathbb{R}$ be the corresponding desired kinematic profiles (knee or ankle joint profiles) as demonstrated by a human. Our goal is a model $f$ which robustly predicts the correct kinematic profiles $y$ from input features $\mathbf{x}$, even under prediction errors in previous time steps, and adapts to new tasks $t$ while preserving performance on older tasks.
To realize such a model, we integrate principles from multitask adaptive learning with a prospective rehearsal approach inspired by model-based RL. Our model builds upon three core features, which we describe in turn: 1) multitask learning via a synthesis of shared learning mechanisms with task-specific adaptability, 2) continual adaptation with task rehearsal in the facet of an evolving architecture to avoid catastrophic forgetting, and 3) a prospective model that imagines and informs about potential deviations to provide robustness against prediction errors.
Figure 2: Summary of the overall approach. For each new task $t$, the data is split into a train $(X^{train}_{t},y^{train}_{t})$ and validation $(X^{val}_{t},y^{val}_{t})$ set. A task-specific layer $f_{t}$ is added to the shared backbone $f_{s}$ and trained to predict the target joint profile $y^{train}_{t}$. Concurrently, a prospective model for the task anticipates potential inputs based on the current joint profiles. The rehearsal data $X^{rehearsal}_{t}$ for all encountered tasks is generated using the prospective model imagined inputs. As new tasks are introduced, both the shared backbone $f_{s}$ and previously established task-specific layers are updated with this rehearsal data to facilitate continual learning.
III-A Multitask learning via generic and task-specific modules
A core challenge in multitask learning is to exploit commonalities between the tasks while maintaining sufficient flexibility to account for the particularities of every single task. Our approach is to compose the gait model $f$ of two functions, namely a shared backbone $f^{s}$ across tasks, followed by task-specific layers $f_{t}$. The overall prediction for an input sample $\{\mathbf{x}_{t,i-T}...\mathbf{x}_{t,i}\}$ for task $t$ at time $i$ is $f(\mathbf{x}_{t,i-T}...\mathbf{x}_{t,i})=f_{t}[f^{s}(\mathbf{x}_{t,i-T}...\mathbf{x}_{t,i})]$. This duality enables the model to distill common features across all tasks and concurrently adapt to the unique aspects of each individual task. We model the shared backbone using temporal convolutions (TCN) [36, 7] due to its efficacy in handling temporal sequences. Task-specific layers are realized using a two-layer feed-forward network, ensuring a lightweight yet effective customization for each task.
III-B Continual adaptation via evolving architecture and rehearsal
In (task) incremental adaptation, the model is trained incrementally with one task at a time. Thus, the model does not require data from all the tasks during training. When a new task $t$ is encountered, a new task-specific prediction layer $f_{t}$ is created and both $f_{t}$ and $f^{s}$ are trained using task-specific data. However, this method is prone to catastrophic forgetting [20, 19], that is, the model forgets the previously trained tasks when it learns a new task. To deal with the forgetting problem, we include samples from prior tasks in the training (i.e., a ’rehearsal’ [37]). However, beyond prior rehearsal approaches, we do not merely copy training data from prior tasks. Instead, we augment it, incorporating prediction errors from the current model $f$ using the following scheme.
III-C Prospective model
A principal challenge in our scenario is that prediction errors at time step $k$ may negatively impact prediction performance at time step $k+1$, such that samples are not identically and independently distributed across time. In more detail, for the human demonstrations, we know that the sensor state $\mathbf{x}_{t,k}$ and desired kinematic profile $y_{t,k}$ is followed by the sensor state $\mathbf{x}_{t,k+1}$. However, if we predict a slightly different kinematic profile $\hat{y}_{t,k}=f(\mathbf{x}_{t,k-T}...\mathbf{x}_{t,k})\neq y_{t,k}$, we will also observe a different sensor state $\hat{\mathbf{x}}_{t,k+1}\neq\mathbf{x}_{t,k+1}$ in the next time step, pushing us away from the training data distribution. Such deviations can accumulate over time, which is a well-known phenomenon in imitation learning [8, 9]. In general, imitation learning is a helpful metaphor for our situation, as we also wish to mimic the demonstration of human experts in a setting where independence over time does not hold. However, past theory on imitation learning has usually assumed constant bounds on the error in every time step. Such constant bounds are not realistic in our scenario with a continuous space. Unfortunately, if we slightly relax the assumption of constant bounds to Lipschitz bounds, the deviation between desired time series and predicted time series can grow exponentially, as we show in the following, simple theorem.
Theorem III.1.
Let $x_{1},\ldots,x_{N}\in\mathcal{X}$ be a time series from space $\mathcal{X}$, equipped with metric $d$. Further, let $f:\mathcal{X}\to\mathcal{Y}$ for some set $\mathcal{Y}$, and let $g:\mathcal{X}\times\mathcal{Y}\to\mathcal{X}$. For any $x^{\prime}_{1}\in\mathcal{X}$, we define the time series $x^{\prime}_{1},\ldots,x^{\prime}_{N}$ via the recursive equation $x^{\prime}_{k+1}=g[x^{\prime}_{k},f(x^{\prime}_{k})]$.
Finally, assume that for all $k\in\{1,\ldots,N-1\}$, for all $x\in\mathcal{X}$, and for some $C\in\mathbb{R}$, the following Lipschitz condition holds: $d\Big{(}g\big{[}x,f(x)\big{]},x_{k+1}\Big{)}\leq C\cdot d(x,x_{k})$. Then, a) we obtain the exponential bound $d(x^{\prime}_{k},x_{k})\leq C^{k-1}\cdot d(x^{\prime}_{1},x_{1})$, and b) this bound is tight, i.e. there exist some time series $x_{1},\ldots,x_{N}$, some $f$ and $g$, as well as some $x^{\prime}_{1}$, such that the bound holds exactly.
Proof.
We obtain a) via induction from $d(x^{\prime}_{k+1},x_{k+1})=d\big{(}g[x^{\prime}_{k},f(x^{\prime}_{k})],x_{k+1}\big{)}\leq C\cdot d(x^{\prime}_{k},x_{k})$. Regarding b), consider the one-dimensional example $\mathcal{X}=\mathcal{Y}=\mathbb{R}$ with metric $d(x,x^{\prime})=|x-x^{\prime}|$, $x_{1}=\ldots=x_{N}=0$, $f(x)=\kappa\cdot x$ for some $\kappa>0$, and $g(x,y)=x+y$. Then, the Lipschitz condition is fulfilled because $d\big{(}g[x,f(x)],x_{k+1}\big{)}=|x+\kappa\cdot x-0|=|1+\kappa|\cdot|x-0|=(1+\kappa)\cdot d(x,x_{k})$, that is, $C=1+\kappa$. Further, for any $x^{\prime}_{1}\in\mathbb{R}$, we obtain $d(x^{\prime}_{k+1},x_{k+1})=|x^{\prime}_{k}+\kappa\cdot x^{\prime}_{k}-0|=(1+\kappa)\cdot|x^{\prime}_{k}-0|=C\cdot d(x^{\prime}_{k},x_{k})$, such that the tight bound follows via induction. ∎
In this theorem, $f$ is our prediction model mapping a sensor state $\mathbf{x}_{k}$ to a target joint profile $y_{k}$, and $g$ describes the dynamics of our system, mapping the current sensor state $\mathbf{x}_{k}$ and joint profile $y_{k}$ to the next sensor state $\mathbf{x}_{k+1}$. The theorem states that, even if our predictive model $f$ is Lipschitz-bounded (i.e., it exactly matches the training data and degrades gracefully beyond the training data), the deviation can grow exponentially over time.
This observation motivates the key innovation of our work, namely a prospective rehearsal. More precisely, we train a prospective model $g$ to mimic the dynamics of the system with $g(\mathbf{x}_{k},{y}_{k})\approx\mathbf{x}_{k+1}$ and imagine the subsequent inputs resulting from the predictions of $f$. We then add data tuples $(\textbf{x}^{\prime}_{k+1},{y}_{k+1})$ with $\textbf{x}^{\prime}_{k+1}=g\big{(}\mathbf{x}_{k},f(\mathbf{x}_{k-T}...\mathbf{x}_{k})\big{)}$ to our training data. These additional training data tuples take the prediction errors of $f$ into account and thus help $f$ to counteract its own prediction errors. In our scenario, ${y}_{k+1}$ is the right target for the input $\textbf{x}^{\prime}_{k+1}$ because ${y}_{k+1}$ represents a target joint profile in a prosthetic limb. This target profile should remain the same in the same gait phase, even if $\textbf{x}^{\prime}_{k+1}$ slightly deviates from $\mathbf{x}_{k+1}$. In a general scenario, other targets might be required. Theoretically speaking, our aim is to achieve a constant bound $d\big{(}g[\mathbf{x}_{k},f(\mathbf{x}_{k-T},...,\mathbf{x}_{k})],\mathbf{x}_{k+1}\big{)}\leq\epsilon$ for small $\epsilon>0$, such that classical imitation learning theory applies [8, 9] and an exponential deviation over time is avoided. We do so by minimizing $d\big{(}g[\mathbf{x}_{k},{y}_{k}],\mathbf{x}_{k+1}\big{)}$ explicitly in Eq. (1).
III-D Summary of proposed approach
Initialize the shared backbone $f^{s}$.
for tasks $t$ from $1,\ldots,L$ do
       Let the data for task $t$ be $(\mathbf{x}_{t,1},{y}_{t,1}),\ldots,(\mathbf{x}_{t,N_{t}},{y}_{t,N_{t}})$.
       Use the first $M_{t}$ steps as training data, the final $N_{t}-M_{t}$ steps as validation data for some $M_{t}<N_{t}$.
       Add a task-specific layer $f_{t}$.
       Train $f^{s}$ and $f_{1},\ldots,f_{t}$ by minimizing the loss:
$\mathcal{L}_{f}=\sum_{k=1}^{M_{t}}\left\lVert{y}_{t,k}-f_{t}[f^{s}(\mathbf{x}_{k},...,\mathbf{x}_{k-T})]\right\rVert^{2}$
$+\sum_{t^{\prime}=1}^{t-1}\sum_{k\in V_{t^{\prime}}}\left\lVert{y}_{t^{\prime},k}-f_{t^{\prime}}[f^{s}(\mathbf{x}_{t^{\prime},k},...,\mathbf{x}_{t^{\prime},k-T})]\right\rVert^{2}$
$+\left\lVert{y}_{t^{\prime},k}-f_{t^{\prime}}[f^{s}(\hat{\mathbf{x}}_{t^{\prime},k},...,\hat{\mathbf{x}}_{t^{\prime},k-T})]\right\rVert^{2}$
Train the prospective model $g_{t}$ for task $t$ by minimizing the loss:
$\mathcal{L}_{g}=\sum_{k=1}^{M_{t}-1}\left\lVert\mathbf{x}_{t,k+1}-g_{t}(\mathbf{x}_{t,k},{y}_{t,k})\right\rVert^{2}$ (1)
      Update the rehearsal data for all tasks.
      for task $t^{\prime}$ from $1,\ldots,t$ do
             $V_{t^{\prime}}\leftarrow$ subsample time steps $k$ from $M_{t^{\prime}},\ldots,N_{t^{\prime}}-1$.
             for time step $k\in V_{t^{\prime}}$ do
$\hat{\mathbf{x}}_{t^{\prime},k}$ $=g_{t^{\prime}}\big{(}\mathbf{x}_{t^{\prime},k-1},$
$f_{t^{\prime}}[f^{s}(\mathbf{x}_{t^{\prime},k-T-1},...,\mathbf{x}_{t^{\prime},k-1})]\big{)}$
$V_{t^{\prime}}\leftarrow(\hat{\mathbf{x}}_{t^{\prime},k},{y}_{t^{\prime},k})$
Algorithm 1 Multitask Prospective Rehearsal-based Adaptive Model
Our overall approach is shown in Algorithm 1 and Figures 1 and 2. We add new tasks $t$ one by one and split the training data for task $t$ into a training and a validation set. We train a newly evolved task-specific layer $f_{t}$ and the shared backbone $f^{s}$ to minimize the distance between $f_{t}[f^{s}(\mathbf{x}_{t,k-T}...\mathbf{x}_{t,k})]$ and ${y}_{t,k}$, and, at the same time, we train the shared backbone $f^{s}$ and the old task layers $f_{t^{\prime}}$ to minimize the distance between $f_{t^{\prime}}[f^{s}(\mathbf{x}_{t^{\prime},k-T}...\mathbf{x}_{t^{\prime},k-1},\hat{\mathbf{x}}_{t^{\prime},k})]$ and ${y}_{t^{\prime},k}$ for the data in the rehearsal buffer. The rehearsal buffer contains both the original validation samples $\mathbf{x}_{t^{\prime},k}$, and the prospective samples $\hat{\mathbf{x}}_{t^{\prime},k}$ for a subsample $V_{t^{\prime}}$ of the time steps $k$ in the validation set. The prospective samples are generated via a task-specific, prospective model $g_{t^{\prime}}$ as $\hat{\mathbf{x}}_{t^{\prime},k}=g_{t^{\prime}}\big{(}\mathbf{x}_{t^{\prime},k-1},f_{t^{\prime}}[f^{s}(\mathbf{x}_{t^{\prime},k-T-1}...\mathbf{x}_{t^{\prime},k-1})]\big{)}$.
IV Experiments
Datasets and metrics. We use three real-world human gait experiments’ datasets. The first two, ENABL3S and Embry [38, 39], are publicly available from popular gait labs, with ENABL3S providing approximately 5.2M training and 1.3M test samples from ten subjects across various locomotion activities, and Embry offering 512K training and 104K test samples from ten subjects also covering a range of movements. The third dataset is a novel collection from our lab, focusing on gait patterns of patients (transtibial amputees), recorded with a 200 Hz infrared motion capture system using twelve cameras, and includes activities like walking, stair ascent, and descent, with around 27K training and 11K testing samples. Ethical clearance for these experiments was obtained from our institutional review board. This dataset, along with the others, includes motion-related variables such as 3D angles of body joints and segments, velocities, and accelerations. To evaluate the accuracy of the desired joint profiles, $y_{k}$, we calculate the coefficient of determination ($R^{2}$), between the ground truth trajectories and model-based joint profiles.
Baselines. We conduct a comprehensive benchmarking of our approach against a range of learning-based limb behavior modeling methods documented in the literature, ranging from traditional linear models and Linear SVR, through advanced nonlinear models like SVR with RBF kernel, Random Forest (RF), and Decision Trees (DT), to neural networks including FFAN and RNNs with LSTM and GRU. Our model outperforms established benchmarks, providing consistent performance across various datasets (Figure 3).
Figure 3: Evaluation of our proposed multitask prospective rehearsal-integrated model across three distinct datasets (ENABL3, Embry et al., Amputees) in comparison with established baselines cited in the literature
Datasets $\rightarrow$ ENABL3S
Embry et. al. 2018
Speed ($ms^{-}1$) / Incline (${}^{\text{o}}$)
Amputees
Tasks $\rightarrow$
Level
Walk
Ramp
Ascent
Ramp
Descent
Stair
Ascent
Stair
Descent
0.8/-10 0.8/0 0.8/10 1.0/-10 1.0/0 1.0/10 1.2/-10 1.2/0 1.2/10
Level
Walk
Stair
Ascent
Stair
Descent
head rehearsal Joint multi-task learning
S conventional 0.83 0.87 0.89 0.91 0.86 0.86 0.88 0.91 0.91 0.94 0.94 0.85 0.93 0.93 0.79 0.79 0.89
prospective (ours) 0.89 0.91 0.92 0.93 0.88 0.88 0.92 0.92 0.93 0.96 0.94 0.89 0.96 0.94 0.93 0.72 0.90
T conventional 0.90 0.94 0.95 0.96 0.93 0.95 0.98 0.94 0.95 0.99 0.96 0.80 0.96 0.98 0.97 0.82 0.92
prospective (ours) 0.92 0.94 0.96 0.96 0.94 0.96 0.98 0.96 0.96 0.99 0.97 0.90 0.97 0.98 0.98 0.91 0.94
head rehearsal Task-incremental learning
S conventional 0.80 0.86 0.84 0.89 0.83 0.76 0.87 0.83 0.81 0.90 0.90 0.74 0.90 0.88 0.94 0.83 0.68
prospective (ours) 0.84 0.90 0.89 0.91 0.89 0.90 0.92 0.92 0.94 0.96 0.95 0.89 0.97 0.95 0.94 0.85 0.72
T conventional 0.88 0.93 0.93 0.94 0.92 0.92 0.98 0.93 0.94 0.98 0.97 0.87 0.97 0.97 0.97 0.91 0.91
prospective (ours) 0.91 0.95 0.95 0.97 0.94 0.96 0.98 0.95 0.97 0.99 0.97 0.91 0.97 0.98 0.98 0.91 0.92
TABLE I: Ablation results: $R^{2}$ values for comparing rehearsal methods and architectures in joint and continual (task-incremental) settings.
Ablations. We assess the contribution of each component on overall performance. We explore both joint and continual learning paradigms, comparing models with shared final layers against those with task-specific final layers. Further, we consider both conventional and prospective rehearsal with each of the shared and task-specific final layer variants. In joint training, the training data is either augmented with the proposed multitask prospective rehearsal from a validation set or simply with the original samples from the validation set itself (conventional rehearsal). Note that, in all conditions, we used the same amount of training data to maintain fairness. Augmentation with prospective rehearsal improved the performance of models with both shared and task-specific final layers. Further, the results provide evidence that task-specific final layers help to account for particularities among tasks. A comparison of performance in task-incremental learning yields similar results. The results demonstrate that the choice of a combination of prospective rehearsal and an evolving architecture, where task-specific final layers are integrated with a shared backbone, yields the best performance across the board (Table I). All following experiments are performed with task-specific final layers.
Figure 4: Comparison of prospective rehearsal with a conventional rehearsal in the face of distribution shift. (Left) Training/test data distributions, their Jenson-Shannon (JS) distances, and $R^{2}$ predictions on four different tasks from the Embry dataset show the prospective rehearsal’s efficacy. (Right) A graph plots the $R^{2}$ difference between methods against JS distance, with a linear trendline indicating that the prospective rehearsal’s benefits increase with greater distribution divergence.
Dataset ENABL3S
Embry et. al. 2018
Speed ($ms^{-}1$)/Incline(${}^{\text{o}}$)
Transtibial
amputees
Task Level Walk Ramp Ascent Ramp Descent Stair Ascent Stair Descent 0.8/-10 0.8/0 0.8/10 1.0/-10 1.0/0 1.0/10 1.2/-10 1.2/0 1.2/10 Level Walk Stair Ascent Stair Descent
SI 0.06 0.45 0.51 0.55 0.93 -0.38 0.33 0.59 0.26 0.40 0.83 0.34 0.66 0.97 0.37 0.31 0.92
EWC 0.13 0.31 0.55 0.52 0.92 0.04 0.33 0.65 0.08 0.54 0.92 0.53 0.77 0.96 -0.30 -0.57 0.85
PNN 0.90 0.91 0.94 0.93 0.91 0.85 0.97 0.96 0.88 0.97 0.96 0.87 0.96 0.98 0.68 0.82 0.79
GEM 0.76 0.91 0.92 0.95 0.93 0.92 0.97 0.93 0.90 0.97 0.96 0.84 0.97 0.97 0.82 0.80 0.86
ER 0.88 0.93 0.93 0.94 0.92 0.92 0.98 0.93 0.94 0.98 0.97 0.87 0.97 0.97 0.97 0.91 0.91
prospective (ours) 0.91 0.95 0.95 0.97 0.94 0.96 0.98 0.95 0.97 0.99 0.97 0.91 0.97 0.98 0.98 0.91 0.92
TABLE II: Comparison of performance ($R^{2}$) of proposed prospective rehearsal with other representative continual learning strategies – SI (Synaptic Intelligence), EWC (Elastic weight consolidation), PNN (Progressive Neural Networks), GEM (Gradient episodic memory), ER (Experience Replay).
Datasets ENABL3S
Embry et. al. 2018
Speed ($ms^{-}1$) / Incline (${}^{\text{o}}$)
Transtibial
Amputees
Tasks
Level
Walk
Ramp
Ascent
Ramp
Descent
Stair
Ascent
Stair
Descent
0.8/-10 0.8/0 0.8/10 1.0/-10 1.0/0 1.0/10 1.2/-10 1.2/0 1.2/-10
Level
Walk
Stair
Ascent
Stair
Descent
Gaussian noise inj. 0.88 0.93 0.93 0.94 0.91 0.90 0.97 0.94 0.91 0.97 0.96 0.81 0.97 0.97 0.92 0.83 0.88
prospective (ours) 0.91 0.95 0.95 0.97 0.94 0.96 0.98 0.95 0.97 0.99 0.97 0.91 0.97 0.98 0.98 0.91 0.92
TABLE III: Comparison of performance ($R^{2}$) of Gaussian noise injection and proposed prospective rehearsal.
Benchmarking continual learning strategies.
In the continual learning paradigm, we compare our approach against regularization strategies like SI and EWC, architectural strategies, PNNs, and state-of-the-art replay-based strategies like GEM and the classic form of experience replay, which we refer to as conventional rehearsal (Table II). We observe that prospective rehearsal performs best across the board. In Table III, we compare data augmentation via Gaussian noise injection to the proposed prospective rehearsal, with prospective rehearsal performing best across the board.
Resilience to distribution shift. In some cases, we do not observe a significant improvement in the performance of the prospective rehearsal over a conventional rehearsal. To explain these results, we inspect the difference in $R^{2}$ between the prospective and conventional rehearsal versus the Jenson-Shannon (JS) distance [40] between training and test distributions (Figure 4). We observe that the error difference positively correlates with JS distance. In other words, the advantage of the prospective rehearsal becomes more pronounced the more training and test distribution deviate. This is in line with our theoretical analysis: The more training and test distribution differ, the more prediction errors of $f$ we expect, which means that our proposed prospective rehearsal helps more. Since real-world prosthesis data is prone to noise and variability from various exteroceptive and interoceptive factors [41], being robust to distribution shifts is an important requirement for this domain, and our approach performs well in this regard.
Figure 5: A) Performance comparison of prospective and conventional rehearsal strategies across different strengths, $\tau$, of adversarial perturbations. B) Performance of linear and non-linear probing on downstream locomotion task classification. Prospective rehearsal training performs best across the board.
Figure 6: (A) (Left) Results of statistical significance tests comparing the performance of conventional and proposed prospective rehearsal for shared and task-specific final layers. (Right) Comparison of performance of the proposed prospective rehearsal with state-of-the-art continual learning strategies. (B) Comparison of the performance of continually trained models with conventional and proposed prospective rehearsal for different levels of input noise. Results are shown for models with task-specific final layers. (C) Plot depicting NRMSE values for test tasks from Embry and ENABL3S datasets post-training, with the X-axis for learned tasks, color for test tasks, and shading indicating standard error across subjects. Statistical significance was assessed using the Wilcoxon-signed rank test with Bonferroni correction, ${}^{****}:p<1e-4,^{***}:p<1e-3,^{**}:p<1e-2$.
Resilience to adversarial perturbations. To further assess the resilience of the prospective rehearsal approach, we compare its outcomes to those without it in the face of adversarial perturbations. These adversarial samples are crafted using the Fast Gradient Sign Method (FGSM) [42] across diverse epsilon thresholds, $\tau$. Notably, our proposed prospective rehearsal technique surpasses the conventional baseline across all evaluated FGSM attack intensities (Figure 5A). Furthermore, as the perturbation intensity escalates, the performance disparity between the prospective and the conventional method widens, reinforcing the superior adaptability of prospective rehearsal under scenarios where test data distribution deviates from the training set. The obtained results are statistically significant.
Probing downstream performance. We evaluate the representation capacity of models trained with prospective rehearsal techniques versus those trained without it, focusing on a downstream challenge—classifying different types of locomotion. We employ both linear and nonlinear (MLP) probes to assess the representation of the shared layers, denoted as $f_{s}(\textbf{x}_{t,k-T}...\textbf{x}_{t,k})$, from models with task-specific heads aimed at classifying the task $t$. The findings indicate that prospective rehearsal yields significantly better task classification performance than its conventional counterpart for both linear and nonlinear probes (Figure 5B). This result underscores that the prospective model-based training confers an added benefit of robustness and adaptability compared to conventional rehearsal techniques.
Resilience to noise. Further, we analyze the robustness of our proposed model against perturbations to the input signals during test time in the form of multivariate white Gaussian noise, $\mathcal{N}(\mathbf{0},\mathbf{\Sigma}\in\mathbb{R}^{d\times d})$. We apply different levels of perturbation $L$, where $L$ is defined as the ratio of the standard deviation of the Gaussian noise to the standard deviation of the test inputs. Figure 6B shows the $R^{2}$ of our proposed rehearsal and a conventional rehearsal for five different tasks on the ENABL3S dataset. While performance generally decreases with higher amounts of noise, the prospective rehearsal yields better $R^{2}$ across the board.
Stability against forgetting. We explore the ability to preserve the knowledge of previously learned tasks after learning new ones. Figure 6C shows how the NRMSE values (lower is better) for each task develop when training subsequent tasks, both for a shared final layer (top) and for a task-specific final layer (bottom). Both architectures preserve the memory of the previous tasks, but the performance on old tasks slightly worsens for a shared final layer, whereas it remains the same for task-specific final layers.
Statistical significance of results. A Wilcoxon-signed rank test revealed that the proposed rehearsal significantly outperforms conventional rehearsal, both for shared as well as task-specific final layers (Figure 6A). Further, task-specific final layers significantly outperform a shared final layer, irrespective of rehearsal strategy. This shows that both the evolving architecture as well as the prospective rehearsal strongly contribute to the improvement in performance using the proposed method (see also Table I). Our experiments also show that the proposed rehearsal significantly outperforms many state-of-the-art continual learning strategies and data augmentation with Gaussian noise (Figure 6A).
V Conclusions
Our research introduces a novel framework for bionic prostheses’ application, adept at handling multiple locomotion tasks, adapting progressively, foreseeing movements, and refining. Central to our method, is a novel prospective rehearsal training scheme that, through empirical studies on diverse datasets, reliably outperforms standard techniques, particularly under challenging conditions of adversaries, distribution shifts, noise, and task transfer. Notably, the prospective rehearsal approach is data-centric, hence, model-agnostic, and can be applicable across various model architectures. While these results are promising, a limitation of our study is the absence of comprehensive clinical trials, largely due to the rigorous and extensive ethical approval process required for such research. Recognizing this, our future work is set to expand upon these findings through more extensive trials, including clinical trials and in-home studies, to validate and refine our model in everyday settings. This adaptable, robust system signals a significant leap forward for prosthetic gait prediction in dynamic settings, offering improved quality of life for individuals with lower limb impairments.